what are the steps for 2a+3b
a=4
b=5

Answer by lenny460(1073)

what are the steps for 2a+3b
a=4
b=5
2a + 3b
2(4) + 3(5)
2 * 4 = 8
3 * 5 = 15
8 + 15 = 23

Answer:
23


obuong3@aol.com